Stan invested $5,000, part at 8% and part at 17%. If the total interest at the end of the year is $490, how much did he invest at 8%?
Answer by checkley71(8403)

.17x+.08(5000-x)=490 or .17x+400-.08x=490 or .17x-.08x=490-400 or
.09x=90 or x=90/.09 or x=1000 thus 1000 was invested @ 17% then 5000-1000=4000
was invested @ 8%
proof .17*1000+.08*4000=490 or 170+320=490 or 490=490